What is the ratio of the shear modulus G to the modulus of elasticity of steel if it has a Poisson’s ratio of 0.25.
1
Expert's answer
2021-07-12T03:36:35-0400


E=2K(1+m)

E/K=2(1+m)

E/K=2(1+0.25)

E/K=2(1.25)

E/K=2.5